OpenStudy (anonymous):

What is a second point on the line passing through the point (2, -3) and having a slope of -1

OpenStudy (anonymous):

(3,-4) maybe, depending on what the first point is... if you are counting it as (2,-3) i think that (3,-4) is the second point.

OpenStudy (anonymous):

Y=mx+b You got points (x,y) = (2,-3) And m= slope which is 1 So -3=1(2) +b B=-5 Sp the line is y=x-5 ... Now pick a number for x X=3 so then y = -2 which gives point (3,-2)